Calculate the standard entropy changes for the following reactions at 25°C:


a) H2(g) + CuO(s) → Cu(s) + H2O(g)

b) 2Al(s) + 3ZnO(s) → Al2O3(s) + 3Zn(s)

c) CH4(g) + 2O2(g) → CO2(g) + 2H2O(l)

d) S(s) + O2(g) → SO2(g)

e) MgCO3(s) → MgO(s) + CO2(g)


1
Expert's answer
2022-06-10T12:17:04-0400

∆S = Sproduced - Sreactant

a.)

H2 + CuO = Cu + H2O

S (H2) = 130,7 J/(mol*K)

S (CuO) = 42,6 J/(mol*K)

S (Cu) = 33,17 J/(mol*K)

S (H2O) = 188,7 J/(mol*K)


∆S = (33,17+188,7) - (130,7+42,6) = 48,57 J/(mol*K)


b.)

2Al + 3ZnO = Al2O3 + 3Zn

S(Al) = 28,27 J/(mol*K)

S(ZnO) = 43,65 J/(mol*K)

S(Al2O3) = 50,92 J/(mol*K)

S(Zn) = 41,72 J/(mol*K)


∆S = (50,92+ 3*41,72) - (2*28,27+3*43,65) = -11,41 J/(mol*K)


c.)

CH4 + 2O2 = CO2 + 2H2O

S(CH4) = 186,3 J/(mol*K)

S(O2) = 205 J/(mol*K)

S(CO2) = 213,79 J/(mol*K)

S(H2O) = 188,7 J/(mol*K)


∆S = (213,79+2*188,7) - (186,3+2*205) = -5,11 J/(mol*K)


d.)

S + O2 = SO2

S(S) = 32,05 J/(mol*K)

S(O2) = 205 J/(mol*K)

S(SO2) = 248,22 J/(mol*K)


∆S = 248,22 - (32,05+205) = 11,17 J/(mol*K)


e.)

MgCO3 = MgO + CO2

S(MgCO3) = 65,84 J/(mol*K)

S(MgO) = 26,9 J/(mol*K)

S(CO2) = 213,79 J/(mol*K)


∆S = 26,9 + 213,79 - 65,84 = 174,84 J/(mol*K)